# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v3

@@ Abstract
Index Terms:
Deep learning, Deep unfolding, Model-driven machine learning, Wireless communication signal processing, Integrated sensing and communication (ISAC), Intelligent reflecting surfaces (IRS), Terahertz, IoT.
@@ I Introduction
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using DNNs presents significant challenges. Data-driven neural network (NN) architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the $tanh$ function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication [12].
@@ I-A Related Works
The growing adoption of deep unfolding in communication systems has led to several surveys covering its diverse applications. Wang et al. [21] reviewed DL and deep unfolding methods for physical layer tasks such as MIMO detection, channel estimation, and decoding. The authors of [22] discussed its use in transceiver design and channel state information (CSI) estimation. In [23], deep learning methods including DNNs, GANs, and recurrent NN (RNN) were reviewed, with deep unfolding applied to Bayesian estimation. The authors of [15] focused on MIMO detection, multiuser precoding, and belief propagation. A broader comparison of model-based, DL, and hybrid methods is found in [24]. Monga et al. [14] focused on algorithm unrolling for signal/image processing. Additional works [25, 13] explored signal processing and optimization. Eldar et al. [12] and Shi et al. [26] addressed DL-based communication frameworks. In [27], the authors highlight the advantages of deep unfolding for physical receiver design, with particular emphasis on its applications in signal detection and channel estimation. A comparison between data-driven and unfolding-based beamforming appears in [28] and Table II.
@@ I-B Contributions
Existing surveys lack coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below:
@@
    This paper presents a systematic, step-by-step approach for developing deep unfolding algorithms, with a detailed tutorial on the alternating direction method of multipliers (ADMM)-based MIMO detection. The tutorial offers insights into the ADMM unfolding process, identifies trainable parameters, and demonstrates the design of the resulting unfolded network.
@@
    We survey the application of deep unfolding across various domains, including signal detection, channel estimation, precoder design, ISAC, decoding, power allocation, and physical layer security. Each section highlights the integration of deep unfolding with key 6G technologies such as IRS, mmWave, and IoT, providing readers with guidance on designing frameworks adaptable to future wireless systems.
@@
    In addition, each section concludes with a comparative analysis table that evaluates prominent unfolded networks across key metrics such as computational complexity, generalization capability, and convergence behavior. These tables highlight the practical trade-offs and deployment boundaries associated with each architecture. Furthermore, we discuss the key takeaways to provide readers with deeper insights into the practical applicability and design considerations of deep unfolding in physical layer communication.
@@
    This paper concludes the survey by outlining a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance.
TABLE II: Existing survey papers on deep unfolding for communication.
SNo. Reference Year Detection Channel Estimation Precoder Design Sensing and Communication Decoding in Error Correcting Codes Power Allocation Security
1 Deep Learning for Wireless Physical Layer: Opportunities and Challenges [21] 2017 $\checkmark$ $\checkmark$ $\checkmark$
2 Model-Driven Deep Learning for Physical Layer Communications [22] 2019 $\checkmark$ $\checkmark$
3 Deep Learning in Physical Layer Communications [23] 2019 $\checkmark$ $\checkmark$
4 Deep Unfolding for Communications Systems: A Survey and Some New Directions [15] 2019 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
5 Wireless Networks Design in the Era of Deep Learning: Model-Based, AI-Based, or Both? [24] 2019 $\checkmark$ $\checkmark$
6 Algorithm Unrolling: Interpretable, Efficient Deep Learning for Signal and Image Processing [14] 2021 $\checkmark$ $\checkmark$
7 Intelligent Radio Signal Processing: A Survey [29] 2021 $\checkmark$ $\checkmark$ $\checkmark$
8 Redefining Wireless Communication for 6G: Signal Processing Meets Deep Learning With Deep Unfolding [25] 2021 $\checkmark$ $\checkmark$ $\checkmark$
9 Model-Based Deep Learning: Key Approaches and Design Guidelines [30] 2021 $\checkmark$
10 Model-Based Deep Learning: On the Intersection of Deep Learning and Optimization [13] 2022 $\checkmark$
11 Machine Learning and Wireless Communications [12] 2022 $\checkmark$ $\checkmark$ $\checkmark$
12 Machine Learning for Large-Scale Optimization in 6G Wireless Networks [26] 2023 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
13 Artificial Intelligence-Empowered Hybrid Multiple-Input/Multiple-Output Beamforming: Learning to Optimize for High-Throughput Scalable MIMO [28] 2024 $\checkmark$
14 Model and Data Dual-Driven: A Promising Receiver Design for 6G Wireless Communications [27] 2025 $\checkmark$ $\checkmark$
15 This Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ I-C Paper Outline
The skeleton of this paper is depicted in Fig.  1.
@@
  • Part A: Section II presents the deep unfolding framework, which combines classical iterative algorithms with DNNs. The understanding of deep unfolding is exemplified via ADMM based MIMO detection. A step-by-step approach is presented for developing the unfolding of ADMM specifically for MIMO detection.
@@
  • Part C : Section X discusses the challenges in adopting the deep unfolding approach in wireless communication and the future scope of research. Section XI concludes the paper.
@@ I-D Notation and Acronyms
[…]
@@ II Deep Unfolding: General Architecture and Fundamentals
[…]
@@
$\underset{\begin{subarray}{c}{\bf{x}}\in{{\cal{X}}}^{2M}\end{subarray}}{\min}$ $\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}$
subject to $\quad\mathbf{x}=\mathbf{z},\mathbf{z}\in{{\cal{X}}}^{2M}.$
@@
where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set $[-1,1]$. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
@@
Furthermore, the layer-wise unfolding of ADMM for MIMO detection is illustrated in Fig. 3. The received symbol vector $\mathbf{y}$ and the channel matrix $\mathbf{H}$ are treated as fixed inputs and are provided to every layer of the network. Each layer of the unfolded network corresponds to one iteration of the ADMM procedure, specifically the update steps in equations (5), (7), and (8). Corresponding to each of these equations, a specific $\rho$ is considered to enable learning. These parameters form the trainable set ${\boldsymbol{\theta}}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$, where each layer $l$ has its own set of penalty coefficients. In the final layer $L$, the network produces the estimate ${\mathbf{z}}^{L}=\hat{\bf{z}}$. For $n=1,\dots,N$, data samples, the output $\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})$ is retrieved from the network as the predicted transmitted symbol for ${\bf{y}}$. The network training loss function formed between the predicted output $\hat{\bf z}$, and ground truth transmitted symbol ${\bf z}_{n}$ is given as:
@@
Figure 4: BER performance comparison for 16$\times$16 MIMO system. For soft projection, $tanh$ is used.
@@
Fig. 4 presents the bit error rate (BER) performance comparison of the unfolded ADMM detector with zero forcing (ZF), minimum mean squared error (MMSE), ADMM and sphere decoder (SD), under QPSK modulation. We consider a $16\times 16$ MIMO system with Rayleigh fading channels. The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and standard ADMM detectors, except for SD. The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters are adaptively trained. This example has been included to provide the readers with an illustrative understanding of the deep unfolding process. The comparative analysis highlights the advantage of the unfolded method over classical approaches. Additional performance improvements can be achieved by optimizing or introducing trainable parameters in the unfolded network.

### v4

@@ I Introduction
Communication problems are inherently influenced by dynamic physical constraints that evolve in real-time. These constraints stem from factors such as the physical environment, hardware limitations, and the specific operational requirements of communication systems. Different communication scenarios impose varying demands on bandwidth, power consumption, and latency, which fluctuate with user density, data rates, and service types. Interpreting and addressing these constraints using DNNs presents significant challenges. Data-driven neural network (NN) architectures were developed primarily to process language and vision data, rendering them suboptimal for wireless communication tasks that involve channels, signals, and outputs governed by complex constraints. Although certain constraints can be handled through the application of specific activation functions, for example, employing the tanhtanh function to enforce a constant modulus constraint in analog beamforming, standard activation functions generally lack the versatility to manage the intricate constraints characteristic of wireless communication systems. To mitigate these challenges, model-driven DL approaches facilitate the development of problem-specific network architectures and customized activation functions, optimized for the unique signal processing requirements of wireless communication [222].
@@ I-A Related Works
The growing application of deep unfolding in communication systems highlights the need for a comprehensive survey in this area. There are limited survey articles that comprehensively discuss deep unfolding applications in the context of the physical layer of communication. Here, we present a few related survey articles on deep unfolding that explore some particular task such as detection, channel estimation, or precoder design. Wang et al. in [199] presented a review of DL and deep unfolding approaches that address physical layer signal processing challenges, including MIMO detection, channel estimation, and channel decoding. The authors of [69] provided an overview of deep unfolding applications in transceiver design, MIMO detection, and channel state information (CSI) estimation and feedback. Similarly, in [161], the authors highlighted the role of DL in communication, with a focus on DNNs, GANs, recurrent neural network (RNN), and end-to-end learning. They also explored deep unfolding for Bayesian optimal estimation, using MIMO detection and channel estimation as case studies. The authors in [14] focused primarily on MIMO detection while exploring its use in multiuser MIMO precoding and belief propagation-based channel decoding. A survey on model-based, DL, and hybrid approaches in the context of wireless communication networks is provided in [227]. In [146], Monga et al. conducted a survey of deep unfolding algorithms specifically implemented for applications in signal and image processing. The authors in  [92] reviewed the implementation of deep unfolding in signal processing tasks. A brief review on deep unfolding in MIMO detection is presented in [177]. The authors in [174] have presented a general framework for optimization that leverages the benefit of deep unfolding with signal detection as one of the case studies. In [222], a chapter is dedicated to model-based machine learning for communication. The deep unfolding networks have been discussed for signal detection, channel estimation, and channel coding. Machine learning-based optimization framework is discussed in [173] with an emphasis on deep unfolding in specific design paradigms such as signal detection. Quoc et al. have presented a survey on application of deep unfolding in signal detection, channel estimation and beamforming in [158]. The authors in [126] explored the application of learning-based optimization techniques to enhance the capacity of communication networks. The survey reviews key domains such as encoding-decoding, equalization, power control, and beamforming, emphasizing the impact of optimization through learning on network performance. A comparative study between purely data-driven and deep unfolding is conducted with complete emphasis on HBF optimization in [176] The significant rise in application of deep unfolding has led to development of theoretical framework to validate and support practical experiments. There are several works [165, 33, 34] which primarily focus on the convergence rate of the unfolded iterative shrinkage-thresholding algorithm (ISTA) and its variants such as learned ISTA (LISTA) [33] and analytic LISTA (ALISTA) [34]. The authors of [35] discuss the motivation and design methodology of deep unfolding, using LISTA as a representative case study. They have studied the design of learned parameter, techniques to design the architecture of the network and parameterize the resulting unfolded network. Certain open mathematical problems pertaining to deep unfolding are also explored in [35].
@@ I-B Contributions
Recently, there has been a surge of application of deep unfolding in various areas of physical layer of wireless communication. The existing survey contributions such as [14, 199, 69], all published before 2020, focus on some particular areas of the physical layer of wireless communication, such as detection, channel estimation, precoder design, or channel coding. In [176], the authors have provided a survey on application of deep unfolding only in the area of hybrid beamforming design. However, the realm of deep unfolding has recently expanded into emerging areas such as sensing and communication, which have not been addressed in any of the previous works. Most of the previous literature discuss the general approach of deep unfolding without dedicated analysis concerning communication perspective. Only a few works such as [146, 173] have given a tutorial description but with a biased emphasis towards signal processing tasks. Therefore, we provide a detailed tutorial describing the deep unfolding in communication perspective and discuss convergence proofs that impact algorithmic stability. In addition, in this work, the application of deep unfolding is explored in the context of critical 6G technologies that are not addressed in existing survey articles such as [14, 199, 69]. Therefore, there is a lack of coverage of recent studies addressing critical 6G technologies, such as mmWave spectrum, new air interface that includes massive MIMO, IRS, next-generation multiple access, and are often limited to specific application areas. Consequently, a comprehensive survey is needed to encompass the broad applications of deep unfolding techniques across various aspects of the physical layer in wireless communication which has not been addressed in existing survey articles [14, 199, 69, 176]. This paper provides an in-depth review of deep unfolding methods applied to the physical layer of wireless communication. The contributions are summarized below:
Figure 1: Typical block diagram of a modern wireless communication system. The section number covering a particular component is shown as well.
@@
    This paper presents a systematic, step-by-step approach for developing deep unfolding algorithms, with a detailed tutorial on the alternating direction method of multipliers (ADMM)-based MIMO detection. The tutorial offers insights into the ADMM unfolding process, identification of the trainable parameters, and demonstrates the design of the resulting unfolded network. Furthermore, we discuss the theoretical foundation of the convergence properties and guidelines for selecting hyperparameters that impact algorithmic stability of deep unfolding. We have highlighted that the convergence properties can be studied under three categories: convergence order, generalization error and normalized mean squared error (NMSE).
@@
    Fig. 1 depicts the principal components of the physical layer in contemporary wireless communication systems. Considering this block diagram, we undertake a systematic survey of each component in details. In particular, we investigate the role of deep unfolding for signal detection, channel estimation, precoder design, integrated sensing and communication (ISAC), decoding, power allocation, and physical-layer security. For each domain, we highlight the integration of deep unfolding with key enabling technologies for sixth-generation (6G) networks, such as intelligent reflecting surfaces (IRS), millimeter-wave (mmWave) communications, and the Internet of Things (IoT). The survey aims to provide a comprehensive perspective on the design of adaptable and forward-looking frameworks for next-generation wireless systems.
@@
    In addition, each section concludes with a cross method comparison table that evaluates prominent unfolded networks across key metrics: computational complexity (floating point operations (FLOPs) and big $\mathcal{O}$ complexity), memory footprint (number of trainable parameters) and convergence speed. These tables highlight the practical trade-offs and deployment boundaries associated with each architecture. Furthermore, we discuss the key takeaways to provide readers with deeper insights into the practical applicability and design considerations of deep unfolding in physical layer communication.
@@
    This paper concludes the survey by outlining a set of challenges in developing deep unfolding models and propose future directions that could guide advancements in deep unfolding techniques to enhance their performance. We point out that challenges in designing a deep unfolding model mainly emerge from scenarios where these models does not offer significant gain in performance-complexity tradeoff. However, we also provide potential solutions that deep unfolding models can benefit from to overcome those challenges.
  5. 5.
    We also provide open source representative works with its github links for different design problems such as detection, channel estimation, ISAC, among others. We believe that this will enable readers to reproduce and develop new unfolding methods.
@@ I-C Paper Outline
The skeleton of this paper is depicted in Fig.  2.
@@
  • Part A: Section II presents the deep unfolding framework, which combines classical iterative algorithms with DNNs. The understanding of deep unfolding is exemplified via ADMM based MIMO detection. A step-by-step approach is presented for developing the unfolding of ADMM specifically for MIMO detection. Also a brief understanding on how trainable parameters impact the convergence order of the deep unfolded network is discussed.
@@
  • Part C : Section X discusses the limitations of existing deep unfolding approaches and the challenges associated with their adoption in wireless communications. We further outline potential directions and solutions to address these limitations, highlighting promising avenues for future research. Section XI concludes the paper.
@@ I-D Notation and Acronyms
Lower case, bold-face lower case, and bold-face upper case letters denote scalars, vectors, and matrices respectively. $(\cdot)^{T}$ and $(\cdot)^{H}$ denote transpose and Hermitian transpose, respectively. $\|\cdot\|$ denotes the Euclidean norm of a vector. $\mathbb{R}^{n}$ and $\mathbb{C}^{n}$ denote $n$-dimensional real and complex vector spaces, respectively. $\mathcal{N}(\mu,\Sigma)$ denotes the multivariate Gaussian distribution with mean vector $\mu$ and covariance matrix $\Sigma$, while $\mathcal{CN}(\mu,\Sigma)$ denotes the complex multi-variate Gaussian distribution. Table I lists the main acronyms used in this paper.
@@ II Deep Unfolding: General Architecture and Fundamentals
$\mathbf{H}\triangleq\begin{bmatrix}\mathfrak{R}(\bar{\mathbf{H}})&-\mathfrak{I}(\bar{\mathbf{H}})\\ \mathfrak{I}(\bar{\mathbf{H}})&\mathfrak{R}(\bar{\mathbf{H}})\end{bmatrix}$,
@@
$\begin{IEEEeqnarraybox}[]{[}][c]{rCl}\underset{{\bf{x}}\in{{\cal{X}}}^{2M}}{\min}&&\quad\quad\left\lVert{\bf{y}}-{\bf{Hx}}\right\rVert_{2}^{2}\\ \text{subject to}&&\quad\quad\mathbf{x}=\mathbf{z},\quad\mathbf{z}\in{{\cal{X}}}^{2M}.\end{IEEEeqnarraybox}$ (3)
@@
where $\Pi_{[-1,1]}({\boldsymbol{a}})$ denotes an element-wise projection of $\boldsymbol{a}$ onto the set [1,1][-1,1]. More specifically, if ${\boldsymbol{b}}=\Pi_{[-1,1]}(\boldsymbol{a})$, we have
@@
Furthermore, the layer-wise unfolding of ADMM for MIMO detection is illustrated in Fig. 4. The received symbol vector $\mathbf{y}$ and the channel matrix $\mathbf{H}$ are treated as fixed inputs and are provided to every layer of the network. Each layer of the unfolded network corresponds to one iteration of the ADMM procedure, specifically the update steps in equations (5), (7), and (8). Corresponding to each of these equations, a specific $\rho$ is considered to enable learning. These parameters form the trainable set ${\boldsymbol{\theta}}=\left\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{1}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{2}^{(l)}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho_{3}^{(l)}}\right\}_{l=1}^{L}$, where each layer $l$ has its own set of penalty coefficients. In the final layer $L$, the network produces the estimate $^{\mathbf{z}}^{L}=\hat{\bf{z}}$. For $n=1,\dots,N$, data samples, the output $\hat{\bf{z}}_{n}({\bf{y}}_{n},{\bf{H}};\rho_{1}^{(l)},\rho_{2}^{(l)},\rho_{3}^{(l)})$ is retrieved from the network as the predicted transmitted symbol for $\bf{y}$}. The network training loss function formed between the predicted output $\hat{\bf z}$, and ground truth transmitted symbol ${\bf z}_{n}$ is given as:
@@
TABLE II: Simulation Parameters.
Parameter Name Value
Number of transmit antennas 16
Number of receive antennas 16
Constellation 4 QAM
Propagation model EVA [49]
Training data size 10000
Testing data size 10000
Epochs 20
Batch size 300
Number of Layers 5
$\Pi_{\text{soft}}$ tanhtanh
@@
Figure 5: BER performance comparison for $\times16$ MIMO system in EVA model [49]. For soft projection, tanhtanh is used.
@@
Fig. 5 presents the bit error rate (BER) performance comparison of the unfolded ADMM detector with zero forcing (ZF), minimum mean squared error (MMSE), ADMM and sphere decoder (SD), under QPSK modulation. For the unfolded ADMM, the penalty parameters and step sizes are considered as learnable parameters. We consider a $16\times 16$ MIMO system in extended vehicular (EVA) propagation model. Table II summarizes the key simulation parameters. The power delay profile of EVA propagation model [49] is given by [0,1.5,1.4,3.6,0.6,9.1,7.0,12.0,16.90,-1.5,-1.4,-3.6,-0.6,-9.1,-7.0,-12.0,-16.9] dB against the excess tap delays [$0,30,150,310,370,710,1090,1730,2510$] ns. The results demonstrate that the unfolded ADMM detector outperforms ZF, MMSE, and classical ADMM detectors. The unfolded ADMM detector performs better due to the use of deep unfolding, where learnable parameters: penalty parameters and step sizes, are adaptively trained. The SD achieves near ML performance and is used often as a benchmark scheme for validating the performance of various detectors. This example has been included to provide the readers with an illustrative understanding of the deep unfolding process. The comparative analysis highlights the advantage of the unfolded method over classical approaches. Additional performance improvements can be achieved by optimizing or introducing trainable parameters in the unfolded network.
The introduction of trainable parameters make it challenging to derive theoretical insights for unfolded algorithms as compared to classical optimization algorithms. For any optimization method, it is crucial to perform convergence analysis. In the literature, the convergence studies of various methods can be classified into three broad categories:
  1. 1.
    Convergence order: In this category, the convergence properties of unfolded networks are explored to provide a comparison with classical methods. Under idealized choices of trainable parameters, different unfolded architectures are studied that yield an accurate solution with a faster convergence order. It is well known that the classical ISTA exhibits sublinear convergence [16]. The work in [165] discusses how learned ISTA (LISTA) or unfolded ISTA produces a sequence of trainable parameters per layer exhibiting a linear convergence order. As discussed in [33], in the unfolded ISTA, due to the transformation of trainable parameters into learned weights, a linear convergence order is achieved. The authors in [35] point out that in deep unfolding, the most essential step in achieving better empirical performance and acceleration is to design the learnable parameters and analyze how the learnable parameters influence each other. For instance, in [33], the parameters related to dictionary based matrices in sparse signal recovery, are not shared across the LISTA network’s layers. The learnable parameters are instead trained in each layer, similar to untying them across the layers. The increased complexity in parameterization enables the network to learn complex mapping and exhibit better acceleration in converging towards the solution. In [34], the authors have proposed analytic LISTA (ALISTA), in which the weight matrices are precomputed through an optimization procedure, thereby eliminating the need for learning them from data. Only a small number of scalar parameters: the layer-wise step size, threshold, and momentum factor are learned. Building upon this framework, the subsequent HyperLISTA model dynamically adapts these parameters in response to the reconstruction error, retraining them for each instance to match the current problem state. This instance adaptive strategy yields a superlinear convergence order, a property that cannot be achieved when a uniform set of optimal parameters is learned across the entire dataset. Although theoretical studies have been attempted on the dependency of parameters on convergence order and stability of the algorithm for ISTA, there is a significant scope to explore theoretical studies on how other classical algorithms such as alternating direction of method of multipliers (ADMM), approximate message passing (AMP), etc., perform compared to their respective unfolded versions. For instance, channel estimation for integrated sensing and communication (ISAC) in intelligent transportation system is proposed in [215]. An unfolded network for channel estimation based on ISTA is designed with trainable parameters, that includes the parameters related to dictionary based matrices in each ISTA iteration and the learnable threshold parameter used in the soft-thresholding operator. In this work, the authors theoretically analyze the proposed unfolded network and establish it achieves a linear convergence order.
  2. 2.
    Generalization error: A fundamental question for any learned model is its reconstruction ability. Especially, the degree to which it can accurately recover a parameter from incomplete or corrupted data. Generalization error or bounds give a probabilistic estimate on the difference between the true error (evaluated over the underlying data distribution) and the empirical error computed from the available training samples for a given hypothesis function. Thus, such bounds predict how well a learned model performs on unseen data and assess whether the training process has led to appropriate generalization [18]. The underlying machine learning algorithm of a learned model can be viewed as stochastic transformation which maps training data to hypotheses. The algorithm stability largely depends on the generalization capability of learning algorithm, and thus is of theoretical interest [23]. As described in [18], based on the training sequence $\mathcal{S}$ and given hypothesis space $\mathcal{H}$, a learning algorithm yields a function $h_{S}\in\mathcal{H}$. The function $h_{S}$ is defined as $h_{S}:\mathcal{X}\rightarrow\mathcal{Y}$ for a training sequence from a distribution $\mathcal{D}$ on $\mathcal{X}\times\mathcal{Y}$. Let us consider a training sequence $\mathcal{S}=\{({\mathbf{x}}_{i},{\mathbf{y}}_{i})\}_{i=1,\dots,m}$ with $m$ i.i.d. samples drawn from the distribution $\mathcal{D}$. Therefore, the empirical loss $\hat{\mathcal{L}}(h)$ of a hypothesis $h$ is the reconstruction error on the training sequence $\mathcal{S}$, as given below:
    $\hat{\mathcal{L}}(h)=\frac{1}{m}\sum_{i=1}^{m}\ell(h,{\mathbf{x}}_{i},{\mathbf{y}}_{i}).$ (10)
    The loss function $\ell$ measures the reconstruction error. Mean squared error (MSE) is the most commonly used loss function. The true loss $\mathcal{L}(h)$ defined as the risk of a hypothesis $h$ is computed by evaluating expectation, $\mathbb{E}[\cdot]$ over a distribution $\mathcal{D}$ as shown below:
    $\mathcal{L}(h)=\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}}[\ell(h,\mathbf{x},\mathbf{y})].$ (11)
    The generalization error of the hypothesis $h_{\mathcal{S}}$ is based on the training samples $\mathcal{S}$. It is defined as the difference between the empirical loss and the true loss as shown below:
    $\text{GE}(h_{\mathcal{S}})=|\hat{\mathcal{L}}(h_{\mathcal{S}})-\mathcal{L}(h_{\mathcal{S}})|.$ (12)
    In [23], the authors have identified several definitions to quantify the stability of a learning algorithm such as: hypothesis stability, point-wise hypothesis stability, error stability, and uniform stability.
    The stability of a learning algorithm is quantified by the bound on the generalization error which must be tight. The trainable parameters in deep unfolding are trained via DL tools so it is imperative that algorithmic stability will be impacted by the choice of these parameters. Thus, the notion of algorithmic stability can offer a theoretical bridge between optimization and learning in deep unfolding. A proper parameterization is fundamental to the empirical performance and acceleration achieved during training of the network. Hence, to maintain convergence and algorithmic stability with robust generalization, the learnable parameters must be carefully designed. Consequently, exploring stability aware parameterization and training strategies presents a promising direction for advancing the theoretical foundation of deep unfolding method.
    Moreover, in [18], the authors have discussed generalization error that relies on Rademacher complexity analysis. The result presented in [18], suggested that deep unfolding models may exhibit better generalization capabilities than purely data driven neural networks, in line with empirical results. For example, the generalization error bound for LISTA depends on the number of layers only logarithmically, whereas generalization error bound for purely data driven neural network scale exponentially in the number of layers. There have been works [178, 104] that discuss generalization error and Rademacher complexity, involving different learnable parameters and a variety of network architectures.
  3. 3.
    NMSE: The convergence rates of different approaches: classical, purely data driven and deep unfolding are examined through the NMSE with respect to the number of layers or iterations required to achieve the optimal solution. Some existing works that use NMSE as a performance metric are as follows, for channel estimation [85, 243, 203], detection [204], precoder design [209], ISAC [94], joint activity detection and channel estimation [172, 250, 139]. Generally, NMSE evaluates the performance of any approach by evaluating the MSE between ground truth, $\mathbf{x}$ and estimated value, $\hat{\mathbf{x}}$ followed by normalization as given below:
[…]
    NMSE is used as a unified performance metric to compare classical iterative algorithms, purely data-driven models, and deep unfolding methods. It provides a direct measure of convergence speed in terms of the number of layers or iterations required to reach a target accuracy. The NMSE-vs-layers plots reported in [172, 85, 243, 250, 203, 209, 139, 204, 94] visually illustrate the convergence behavior of unfolded networks relative to benchmark algorithms. These results consistently show that NMSE decreases as the number of layers or iterations increases. Classical iterative methods exhibit slow NMSE reduction because they typically require many iterations to converge. In contrast, deep unfolding achieves the same NMSE performance with significantly fewer layers, outperforming both purely data-driven models and classical approaches [85, 172].
    While NMSE provides a direct measure of reconstruction accuracy with respect to the number of layers, it does not capture characteristics such as the convergence order of the algorithm or its susceptibility to overfitting. These aspects are instead characterized by convergence order analysis and generalization error, which together describe the theoretical convergence behavior and algorithmic stability of deep unfolding methods. However, NMSE can be complemented with these analyses to obtain an understanding of the efficiency and the reliability of unfolded architectures.
The existing theoretical studies have examined the convergence order [165, 33, 34] and generalization error [18, 23, 178] of unfolded networks. However, the authors in [167] have provided optimization guarantees showing that unfolded networks can achieve near-zero training loss as the number of learning epochs increases. Their analysis focuses on how the training loss decreases with more epochs while keeping the number of layers fixed. In their numerical results, they show that purely data-driven networks require significantly more parameters to achieve the same near-zero training loss compared to LISTA and ADMM-based compressive sensing networks. This highlights the parameter efficiency and training advantages of unfolded architectures.